How many electrons can be assigned to orbitals designated by the quantum numbers n = 2, ℓ = 1?
1
Expert's answer
2020-10-14T05:59:46-0400

n=2,l=1, represent the 2p orbital,which contain maximum 6 electron, p subshell has three orbital according to (2l+1) rule, here l =1, so (2l+1)=2*1+1=3,each orbital contain two electrons,hence three orbital contain (3*2)=6 electrons